OpenStudy (zeig_101):

Real quick help please: What is the sum of the geometric series 15Σx=0 2(1/4)^x rounded to the nearest whole number?

OpenStudy (kirbykirby):

\[ 15 \sum_{x=0}^2\left(\frac{1}{4}\right)^x \,\, \text{?}\]

OpenStudy (zeig_101):

Yes, with a 2 in front of 1/4

OpenStudy (zeig_101):

and the 15 is on top of the sigma

OpenStudy (kirbykirby):

\[\sum_{x=0}^{15}2\left(\frac{1}{4}\right)^x \]

OpenStudy (zeig_101):

yes, that is perfect.

OpenStudy (zeig_101):

My options are 1, 2, 3, and 4.

OpenStudy (kirbykirby):

It's 3 to the nearest integer

OpenStudy (zeig_101):

It is?

OpenStudy (zeig_101):

Can you show me how to figure that out?

OpenStudy (kirbykirby):

There is a formula for geometric series: \[ \large \sum_{x=0}^n ar^x=a\frac{1-r^{n+1}}{1-r}\]. Here, \(a\) is any number (n your example it is 2), and \(r\) must be between 0 and 1, and here \(r=1/4\)

OpenStudy (kirbykirby):

then you calculate \[2\cdot \frac{1-(\frac{1}{4})^{16}}{1-\frac{1}{4}} \]

OpenStudy (zeig_101):

okay Where does the 16th power on 1/4 come from?

OpenStudy (kirbykirby):

Your summation goes from 0 to 15. The formula says you have "0 to n", so n=15. But look in the formula you have \(1-r^{n+1}\), so you have \(1-r^{15+1}\)

OpenStudy (kirbykirby):

Oh and I said r should be between 0 and 1, but actually it's valid for |r| <1